OpenStudy (anonymous):

A square has sides of length (3x+5) what is the Perimeter and Area of this square?

OpenStudy (anonymous):

Area of square = s*s Perimeter of square = 4s

OpenStudy (anonymous):

Perimeter = 4(3x+5) Area = (3x+5)(3x+5)

OpenStudy (anonymous):

so the perimeter would be 12x+20?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and the Area 9x^2+30x+25?

OpenStudy (anonymous):

yes...square units for area just units for perimeter

OpenStudy (anonymous):

ok thank you :)